# Instance-dependent uniform tail bounds for empirical processes

### 07 Feb 2024 v4

@@ 1 Introduction
Section 2 provides a more precise statement of the problem of interest. In Section Section 3 we consider two illustrative examples involving Gaussian processes. In particular, in Section 3.1, we discuss the problem studied by [LM23] in more details. Furthermore, in Section 3.2 we use instance-dependent bounds to formulate confidence intervals for the order statistics of a general Gaussian vector. The general instance-dependent tail bounds are presented in Section 4. Our calculations in Section 4.3 to derive more explicit expressions in the case of function classes in (exponential type) Orlicz spaces, can be of independent interest.
@@ 2 Preliminaries and Problem Setup
Let $\mathcal{F}$ denote a finite but arbitrarily large subset [^1] In many situations infinite function classes can be considered as well, but a completely rigorous analysis for the problems of interest requires the measurability issues to be addressed, e.g., as in [Pol84, Appx. C]. of a vector space $\mathbb{V}$ of centered functions from $\mathcal{X}$ to $\mathbb{R}$ whose cumulant generatig function is finite in a neighborhood of the origin. Specifically, for every $f\in\mathbb{V}$ we have
@@
The derivations leading to (6) are not intended to produce the most refined bounds. In fact, it is known that the left-hand side of (6) is of order $\sqrt{\log\log(n)/n}$ (see, e.g., [Csa77], [GK06, Example 2.7], and [WS09, Chapter 16]).
@@ 3.1 Marginals of a Gaussian Vector
3.2 Confidence Intervals for Gaussian Order Statistics
@@
In this subsection we derive confidence intervals for the order statistics of general Gaussian random vectors, as another example where instance-dependent tail bounds can be applied. Similar to Proposition 1, we use the Gaussian concentration inequality, as well as a bound on the expected supremum of canonical Gaussian processes over (symmetric) polytopes [Tal14, Theorem 2.4.18] as mentioned in [vHan18a, Section 3.3]. These tools allow us to express the upper and lower bounds of the confidence interval in more explicit terms.
Example 2.
@@
Our goal is to find an upper and lower bounds for the $m$-th largest entry of a parameter vector $\theta\in\mathbb{R}^{d}$ in magnitude. We are only given $\widehat{\theta}=\theta+X$, where $X$ is a zero-mean Gaussian random variable with covariance $\Sigma=\operatorname{\mathbb{E}}XX^{\scriptscriptstyle\textup{{T}}}$. For any vector $v$ we denote by $|v|$ and $v^{\downarrow}$, respectively, the vector of the absolute values of the entries of $v$, and the vector of the entries of $v$ sorted in decreasing order. Therefore, the $m$-th largest magnitude of the entries of a vector $v$ can be expressed as $|v|^{\downarrow}_{m}$. Furthermore, for any subset $S$ of $[d]\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\{1,\dotsc,d\}$ let $v_{S}$ denote the vector that is identical to $v$ at the indices in $S$, and zero at the other indices. Expressing $|v|^{\downarrow}_{m}$ in an equivalent “min-max” form, we can write
@@
$|\theta|^{\downarrow}_{m}$ $=\min_{S\in\binom{[d]}{d-m+1}}\max_{{\left\lVert u\right\rVert}_{1}\leq 1}{\langle u,\theta_{S}\rangle}$
@@
$=\min_{S\in\binom{[d]}{d-m+1}}\max_{{\left\lVert u\right\rVert}_{1}\leq 1}{\langle u,\widehat{\theta}_{S}\rangle}-{\langle u,X_{S}\rangle}\,,$ (7)
where $\binom{[d]}{d-m+1}$ is the set of subsets of $[d]$ of size $d-m+1$. We use the shorthand $\Sigma_{S}=\operatorname{\mathbb{E}}X_{S}X_{S}^{\scriptscriptstyle\textup{{T}}}$, which is the same as $\Sigma$ except for the rows and columns not in $S$, where it has all zero entries. For a nonnegative integer $k\leq r$, let $\pi_{S}$ and $\pi_{S}^{\perp}$, respectively, denote the orthogonal projections onto the span of the top $k$ eigenvectors of $\Sigma_{S}$, and its orthogonal complement. Then, with $G\sim\mathrm{Normal}(0,I)$ we can write
${\langle u,\pi X_{S}\rangle}$ =$\overset{\mathrm{d}}{=}{\langle\Sigma_{S}^{1/2}\pi_{S}u,\pi_{S}G\rangle}$
@@
$={\left\lVert u\right\rVert}_{\pi_{S}\Sigma_{S}\pi_{S}}{\left\lVert\pi_{S}G\right\rVert}_{2}$
@@
thereby, on the same event, for all $u\in\mathbb{R}^{d}$ we have
@@
Furthermore, let us define the vector $\sigma=\sigma(S,k)$ such that $\sigma_{i}=(\pi_{S}^{\perp}\Sigma_{S}\pi_{S}^{\perp})^{1/2}_{i,i}$ for $i\in[d]$. Then, defining
$\sigma^{*}(S,k)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\max_{i\in[d]}\sigma^{\downarrow}_{i}\sqrt{\log(i+1)}\,,$
with probability at least $1-e^{-r}$ we have
${\langle u,\pi^{\perp}X_{S}\rangle}$ $\leq{\left\lVert u\right\rVert}_{1}{\left\lVert\pi_{S}^{\perp}X_{S}\right\rVert}_{\infty}$
@@
where the second line follows from the Gaussian concentration inequality, and the third line follows from [Tal14, Theorem 2.4.18] (and the subsequent discussion) for some absolute constant $C>0$. Adding the derived inequalities, with probability at least $1-2e^{-r}$, for all $u\in\mathbb{R}^{d}$, we can guarantee
@@
$|\theta|_{m}^{\downarrow}$ $\geq\max_{{\left\lVert u\right\rVert}_{1}\leq 1}{\langle u,\widehat{\theta}_{S}\rangle}-{\left\lVert u\right\rVert}_{1}(C\sigma^{*}(S,k)+\sqrt{2r}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})-{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+k)}\,,$
@@
$|\theta|_{m}^{\downarrow}$ $\leq\max_{{\left\lVert u\right\rVert}_{1}\leq 1}{\langle u,\widehat{\theta}_{S}\rangle}+{\left\lVert u\right\rVert}_{1}(C\sigma^{*}(S,k)+\sqrt{2r}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})+{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+k)}\,.$
We can choose a specific $S$ independent of $\widehat{\theta}$, e.g., the indices of the $d-m+1$ smallest entries of $|\theta|$, or apply the union bound and choose the $S\in\binom{[d]}{d-m+1}$ that minimize/maximize the above upper/lower bound, at the cost of increasing $r$ by $\log\binom{d}{d-m+1}\asymp\min\{d-m+1,m-1\}\log d$.
For comparison, if we choose $k=0$, the left-hand side of (8) vanishes. Therefore, the inequality (9) can be replaced by
${\langle u,X_{S}\rangle}$ $\leq{\left\lVert u\right\rVert}_{1}(C\sigma^{*}(S,0)+\sqrt{2r}{\left\lVert\sigma(S,0)\right\rVert}_{\infty})\,,$
which leads to the bound
@@
$||\theta|_{m}^{\downarrow}-{\left\lVert\widehat{\theta}_{S}\right\rVert}_{\infty}|$ $\leq C\sigma^{*}(S,0)+\sqrt{2r}{\left\lVert\sigma(S,0)\right\rVert}_{\infty}\,.$

### 21 Mar 2024 v5

@@ 1 Introduction
Section 2 provides a more precise statement of the problem of interest. In Section Section 3 we consider two illustrative examples involving Gaussian processes. In particular, in Section 3.1, we discuss the problem studied by [LM23] in more details. Furthermore, in Section 3.2 we use instance-dependent bounds to formulate confidence intervals for the $m$th largest mean of a general Gaussian vector. The general instance-dependent tail bounds are presented in Section 4. Our calculations in Section 4.3 to derive more explicit expressions in the case of function classes in (exponential type) Orlicz spaces, can be of independent interest.
@@ 2 Preliminaries and Problem Setup
Let $\mathcal{F}$ denote a finite but arbitrarily large subset [^1] In many situations infinite function classes can be considered as well, but a completely rigorous analysis for the problems of interest requires the measurability issues to be addressed, e.g., as in [Pol84, Appx. C]. of a vector space $\mathbb{V}$ of centered functions from $\mathcal{X}$ to $\mathbb{R}$ whose cumulant generating function is finite in a neighborhood of the origin. Specifically, for every $f\in\mathbb{V}$ we have
@@
The derivations leading to (6) are not intended to produce the most refined bounds. In fact, it is known that the left-hand side of (6) is of order $\sqrt{\log\log(n)/n}$ (see, e.g., [Csá77], [GK06, Example 2.7], and [WS09, Chapter 16]).
@@ 3.1 Marginals of a Gaussian Vector
3.2 Confidence Intervals for the “Middle-Ranked” Means of Correlated Gaussians
@@
In this subsection we derive confidence intervals for the $m$-th largest mean of correlated Gaussian random variables, as another example where instance-dependent tail bounds can be applied. Similar to Proposition 1, we use the Gaussian concentration inequality, as well as a bound on the expected supremum of canonical Gaussian processes over (symmetric) polytopes [Tal14, Proposition 2.4.16 and Theorem 2.4.18] (see also the discussion in [vHan18, Section 3.3]). These tools allow us to express the upper and lower bounds of the confidence interval in more explicit terms.
@@
Our goal is to find an upper and lower bounds for the $m$-th largest entry of a parameter vector $\theta\in\mathbb{R}^{d}$ for $m=o(d)$. We are only given $\widehat{\theta}=\theta+X$, where $X$ is a zero-mean Gaussian random variable with covariance $\Sigma=\operatorname{\mathbb{E}}XX^{\scriptscriptstyle\textup{{T}}}$. We assume that $\Sigma$ is known, and, without loss of generality, it is full-rank. For any vector $v$ we denote by $v^{\downarrow}$ the vector of the entries of $v$ sorted in decreasing order. Therefore, the $m$-th largest entry of a vector $v$ can be expressed as $v^{\downarrow}_{m}$. Furthermore, for any subset $S$ of $[d]\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\{1,\dotsc,d\}$ let $v_{S}\in\mathbb{R}^{|S|}$ denote the restriction of $v$ to the entries indexed by $S$. We also use the shorthand $\Sigma_{S}=\operatorname{\mathbb{E}}X_{S}X_{S}^{\scriptscriptstyle\textup{{T}}}$, which is the same as $\Sigma$ restricted to the rows and columns in $S$. By $\binom{[d]}{\ell}$, we denote the set of subsets of $[d]$ of size $\ell$, and we write $\triangle^{\ell}$ to denote the unit simplex in $\mathbb{R}^{\ell}$
Perhaps the simplest approach for our problem is to use the inequality
@@
$\left|\theta^{\downarrow}_{m}-\widehat{\theta}^{\downarrow}_{m}\right|\leq{\left\lVert\theta-\widehat{\theta}\right\rVert}_{\infty}={\left\lVert X\right\rVert}_{\infty}\,,$
that suggests a confidence interval centered at the plug-in estimator $\widehat{\theta}^{\downarrow}_{m}$ whose width is no less than $2{\left\lVert X\right\rVert}_{\infty}$. The Gaussian concentration inequality then guarantees that ${\left\lVert X\right\rVert}_{\infty}\leq\operatorname{\mathbb{E}}{\left\lVert X\right\rVert}_{\infty}+\sqrt{2r}\max_{i\in[d]}\Sigma^{1/2}_{i,i}$ with probability at least $1-e^{-r}$. Furthermore, we can bound $\operatorname{\mathbb{E}}{\left\lVert X\right\rVert}_{\infty}$, viewed as the expected supremum of a canonical Gaussian process over a (symmetric) polytope, using [Tal14, Proposition 2.4.16 and Theorem 2.4.18]. Denoting the $i$th largest diagonal entry of $\Sigma$ by $\Sigma^{\downarrow}_{i,i}$, for some constant $C>0$ we have
$\left|\theta^{\downarrow}_{m}-\widehat{\theta}^{\downarrow}_{m}\right|$ $\leq C\max_{i\in[d]}\sqrt{\Sigma^{\downarrow}_{i,i}\log(i+1)}+\sqrt{2r}\sqrt{\Sigma^{\downarrow}_{1,1}}\,.$ (7)
Using the instance-dependent uniform tail bounds, we establish a more refined confidence interval for $\theta^{\downarrow}_{m}$ in the following proposition.
Proposition 2.
Let $m=o(d)$ be a positive integer, $r\in\mathbb{R}_{\geq 0}$, and $k\leq\min\{r,m\}$ be a nonnegative integer. Furthermore, for any nonempty set $S\subseteq[d]$ denote by $\Sigma_{S,k}$ the best rank-$k$ approximation of $\Sigma_{S}$ with respect to the operator norm, and define the vector $\sigma=\sigma(S,k)$ such that $\sigma_{i}=\sqrt{(\Sigma_{S}-\Sigma_{S,k})_{i,i}}$ for $i\in[|S|]$. Then, defining
$\sigma^{*}(S,k)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\max_{i\in[|S|]}\sigma^{\downarrow}_{i}\sqrt{\log(i+1)}\,,$
@@
$Q_{S,\beta}(\vartheta)$ $=\max_{u\in\triangle^{|S|}}{\langle u,\vartheta\rangle}-\beta{\left\lVert u\right\rVert}_{\Sigma_{S}}\,,$
which implicitly depends on $\Sigma_{S}$, and
$\beta_{r,m,k}$ $=\sqrt{r+m+m\log(d/m)+k}\,,$
with probability at least $1-4e^{-r}$ we have
$\theta^{\downarrow}_{m}$ $\geq\min_{S\in\binom{[d]}{d-m+1}}Q_{S,\beta_{r,m,k}}(\widehat{\theta}_{S})-C\sigma^{*}(S,k)-\sqrt{2}\beta_{r,m,k}{\left\lVert\sigma(S,k)\right\rVert}_{\infty}\,,$ (8)
and
$\theta^{\downarrow}_{m}$ $\leq\max_{S\in\binom{[d]}{m}}-Q_{S,\beta_{r,m,k}}(-\widehat{\theta}_{S})+C\sigma^{*}(S,k)+\sqrt{2}\beta_{r,m,k}{\left\lVert\sigma(S,k)\right\rVert}_{\infty}\,.$ (9)
Proof.
We express $\theta^{\downarrow}_{m}$ in an equivalent min-max variational form as
$\theta^{\downarrow}_{m}$ $=\min_{S\in\binom{[d]}{d-m+1}}\max_{u\in\triangle^{d-m+1}}{\langle u,\theta_{S}\rangle}$
$=\min_{S\in\binom{[d]}{d-m+1}}\max_{u\in\triangle^{d-m+1}}{\langle u,\widehat{\theta}_{S}\rangle}-{\langle u,X_{S}\rangle}\,.$ (10)
For the prescribed nonnegative integer $k\leq r$, let $\pi_{S}$ and $\pi_{S}^{\perp}$, respectively, denote the orthogonal projections onto the range and the nullspace of $\Sigma_{S,k}$. Then, with $G\sim\mathrm{Normal}(0,I)$ we can write
${\langle u,\pi_{S}X_{S}\rangle}$ =$\overset{\mathrm{d.}}{=}{\langle\Sigma_{S}^{1/2}\pi_{S}u,\pi_{S}G\rangle}$
@@
$={\left\lVert u\right\rVert}_{\Sigma_{S,k}}{\left\lVert\pi_{S}G\right\rVert}_{2}$
@@
thereby, on the same event, for all $u\in\mathbb{R}^{d-m+1}$ we have
@@
Furthermore, recalling the definition of $\sigma=\sigma(S,k)$, with probability at least $1-e^{-r}$ we have
${\langle u,\pi^{\perp}_{S}X_{S}\rangle}$ $\leq{\left\lVert u\right\rVert}_{1}{\left\lVert\pi_{S}^{\perp}X_{S}\right\rVert}_{\infty}$
@@
where the second line follows from the Gaussian concentration inequality, and the third line follows from [Tal14, Proposition 2.4.16 and the remarks after Theorem 2.4.18] for some absolute constant $C>0$. Adding the derived inequalities, with probability at least $1-2e^{-r}$, for all $u\in\mathbb{R}^{d-m+1}$, we can guarantee
@@
$\max_{u\in\triangle^{d-m+1}}{\langle u,\theta_{S}\rangle}$ $\geq\max_{u\in\triangle^{d-m+1}}{\langle u,\widehat{\theta}_{S}\rangle}-(C\sigma^{*}(S,k)+\sqrt{2r}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})-{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+k)}\,.$
To obtain a lower bound for $\theta^{\downarrow}_{m}$, we can choose $S$ to be the indices of the $d-m+1$ smallest entries of $\theta$. But to be truly agnostic to the choice of $\theta$, we need to invoke the union bound and minimize the lower bound with respect to $S\in\binom{[d]}{d-m+1}$, at the cost of increasing $r$ by $m+m\log(d/m)>\log\tbinom{d}{m-1}$. The resulting inequality is then
$\theta^{\downarrow}_{m}$ $\geq\min_{S\in\binom{[d]}{d-m+1}}\max_{u\in\triangle^{d-m+1}}\Big{(}{\langle u,\widehat{\theta}_{S}\rangle}-(C\sigma^{*}(S,k)+\sqrt{2(r+m+m\log(d/m))}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})$
$-{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+m+m\log(d/m)+k)}\Big{)}\,,$
which, by identifying the expressions of $\beta_{r,m,k}$ and $Q_{S,\beta_{r,m,k}}(\cdot)$, is equivalent to (8). To establish the upper bound (9), observe that $\theta^{\downarrow}_{m}=-(-\theta)^{\downarrow}_{d-m+1}$, which allows us to reuse the inequalities above to derive an upper bounds for $\theta^{\downarrow}_{m}$ through the lower bound for $(-\theta)^{\downarrow}_{d-m+1}$. ∎
It is worth mentioning that for $k=0$, the left-hand side of (11) vanishes, thereby we can improve the inequality (12) to
${\langle u,X_{S}\rangle}$ $\leq{\left\lVert u\right\rVert}_{1}(C\sigma^{*}(S,0)+\sqrt{2r}{\left\lVert\sigma(S,0)\right\rVert}_{\infty})\,.$
Consequently, for $k=0$ the corresponding bounds are in fact
$\theta^{\downarrow}_{m}$ $\geq\min_{S\in\binom{[d]}{d-m+1}}\left(\max_{i\in S}\widehat{\theta}_{i}-C\sigma^{*}(S,0)-\sqrt{2(r+m+m\log(d/m))}{\left\lVert\sigma(S,0)\right\rVert}_{\infty}\right)\,,$
@@
$\theta^{\downarrow}_{m}$ $\leq\max_{S\in\binom{[d]}{m}}\left(\min_{i\in S}\widehat{\theta}_{i}+C\sigma^{*}(S,0)+\sqrt{2(r+m+m\log(d/m))}{\left\lVert\sigma(S,0)\right\rVert}_{\infty}\right)\,.$
If in addition to the assumption $m=o(d)$, we also have $m\lesssim r$ (i.e., $m\leq cr$ for some fixed constant $c>0$), the bounds above reproduce (7) up to an extra logarithmic factor for the term $\sqrt{\Sigma^{\downarrow}_{1,1}}$.
We also have the following minimax lower bounds for estimating $\theta^{\downarrow}_{m}$.
Proposition 3.
For $\kappa\geq 0$, let $\Theta=\{\theta\in\mathbb{R}^{d}\colon{\left\lVert\Sigma^{-1/2}\theta\right\rVert}_{2}\leq\kappa\}$ be a compact domain of parameters. With $\delta_{m}\geq 0$ defined as
$\delta_{m}$ $=\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}+\sup_{\eta\in\Theta}\eta^{\downarrow}_{d-m+1}\,.$
For any estimator $g(\widehat{\theta})$ of $\theta^{\downarrow}_{m}$ we have
$\sup_{\theta\in\Theta}\operatorname{\mathbb{E}}(g(\widehat{\theta})-\theta^{\downarrow}_{m})^{2}$ $\geq\frac{\delta_{m}^{2}}{8e\max\{1,2\kappa^{2}\}}\,.$ (13)
Furthermore, we have
$\sup_{\theta\in\Theta}\operatorname{\mathbb{P}}\left(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|>\frac{\delta_{m}}{3\max\{1,\sqrt{2}\kappa\}}\right)$ $\geq\frac{1}{2e}\,.$
Proof.
Le Cam’s two point method [PW24, Theorem 31.1] (see also [Yu97, Lemma 1]) guarantees that
$\sup_{\theta\in\Theta}\operatorname{\mathbb{E}}(g(\widehat{\theta})-\theta^{\downarrow}_{m})^{2}$ $\geq\sup_{\theta,\eta\in\Theta}\frac{1}{4}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}\left(1-D_{\mathrm{TV}}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{\eta})\right)\,,$
where $D_{\mathrm{TV}}(\cdot,\cdot)$ denotes the total variation distance, and $\operatorname{\mathbb{P}}_{\theta}=\mathrm{Normal}(\theta,\Sigma)$ and $\operatorname{\mathbb{P}}_{\eta}=\mathrm{Normal}(\eta,\Sigma)$. The “simplified” Bretagnolle–Huber inequality [Tsy08, Equation 2.25] (see also [Can23] for a broader context) guarantees that
$D_{\mathrm{TV}}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{\eta})$ $\leq 1-\frac{1}{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}\,,$ (14)
using which we obtain
$\sup_{\theta\in\Theta}\operatorname{\mathbb{E}}(g(\widehat{\theta})-\theta^{\downarrow}_{m})^{2}$ $\geq\sup_{\theta,\eta\in\Theta}\frac{1}{8}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}$
$=\sup_{\theta,\eta\in\Theta}\sup_{b\in[0,1]}\frac{1}{8}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}be^{-b(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}$
$\geq\frac{1}{8e\max\{1,2\kappa^{2}\}}\sup_{\theta,\eta\in\Theta}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}\,,$
where the second line follows from the fact that $\Theta$ is star-shaped, and the third line follows from the inequality $\max_{b\in[0,1]}be^{-bz}\geq e^{-1}/\max\{1,z\}$ for $z\geq 0$. We can derive (13) from this lower bound using the fact that
$\sup_{\theta,\eta\in\Theta}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}$ $=\left(\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}-\inf_{\eta\in\Theta}\eta^{\downarrow}_{m}\right)^{2}$
$=\left(\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}+\sup_{\eta\in\Theta}\eta^{\downarrow}_{d-m+1}\right)^{2}\,,$
where the latter equation follows from the symmetry of the set $\Theta$, and the fact that $-\eta^{\downarrow}_{m}=(-\eta)^{\downarrow}_{d-m+1}$.
Furthermore, it follows from the definition of the total variation distance and (14) that for any $c\geq 0$ we have
$|\operatorname{\mathbb{P}}(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|>c)-\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\theta^{\downarrow}_{m}|>c)|$ $\leq 1-\frac{1}{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}\,.$
In particular, for any $c\leq|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|/2$, together with the inequality
$\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\theta^{\downarrow}_{m}|>\frac{1}{2}|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|)$ $\geq\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\eta^{\downarrow}_{m}|<\frac{1}{2}|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|)\,,$
which follows from the triangle inequality, we obtain
$\operatorname{\mathbb{P}}(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|\geq c)+\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\eta^{\downarrow}_{m}|\geq c)$ $\geq\frac{1}{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}\,.$
Therefore, if there exists a pair $\theta,\eta\in\Theta$ such that $(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)\leq 2$ and $|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|/2\geq c$, then
$\sup_{\theta\in\Theta}\operatorname{\mathbb{P}}(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|\geq c)$ $\geq\frac{1}{2e}\,.$
The desired result follows by setting $c=\sup_{\theta,\eta\in\Theta}|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|/(3\max\{1,\sqrt{2}\kappa\})$ which meets the required conditions. ∎
Because of the complicated and implicit form of the expressions in (8) and (9), it is difficult to compare—in full generality—the width of the confidence interval provided by Proposition 2 and the minimax lower bound of Proposition 3. We only focus on the special case where $\Sigma$ is diagonal. Furthermore, for the sake of simpler calculations we use the lower bound
$Q_{S,\beta}(\widehat{\theta}_{S})$ $\geq\max_{i\in S}\left(\widehat{\theta}_{i}-\beta\sqrt{\Sigma_{i,i}}\right)\,.$
The width of the confidence interval expressed by (8) and (9), which we denote by $\Delta_{m}$, can be bounded as
$\Delta_{m}=\max_{S\in\binom{[d]}{m}\,,S^{\prime}\in\binom{[d]}{d-m+1}}$ $\Big{(}-Q_{S,\beta_{r,m,k}}(-\widehat{\theta}_{S})-Q_{S^{\prime},\beta_{r,m,k}}(\widehat{\theta}_{S^{\prime}})$
$+C(\sigma^{*}(S,k)+\sigma^{*}(S^{\prime},k))$
$+\sqrt{2}\beta_{r,m,k}({\left\lVert\sigma(S,k)\right\rVert}_{\infty}+{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty})\Big{)}$
$\leq\max_{S\in\binom{[d]}{m}\,,S^{\prime}\in\binom{[d]}{d-m+1}}\Big{(}\min_{i\in S\,,j\in S^{\prime}}$ $\widehat{\theta}_{i}-\widehat{\theta}_{j}+\beta_{r,m,k}(\sqrt{\Sigma_{i,i}}+\sqrt{\Sigma_{j,j}})$
$+C(\sigma^{*}(S,k)+\sigma^{*}(S^{\prime},k))$
$+\sqrt{2}\beta_{r,m,k}({\left\lVert\sigma(S,k)\right\rVert}_{\infty}+{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty})\Big{)}$
$\leq\max_{S\in\binom{[d]}{m}\,,S^{\prime}\in\binom{[d]}{d-m+1}}\Big{(}\min_{i\in S\cap S^{\prime}}$ $2\beta_{r,m,k}\sqrt{\Sigma_{i,i}}+C(\sigma^{*}(S,k)+\sigma^{*}(S^{\prime},k))$
$+\sqrt{2}\beta_{r,m,k}({\left\lVert\sigma(S,k)\right\rVert}_{\infty}+{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty})\Big{)}\,,$
where the second inequality holds because $|S\cap S^{\prime}|=|S|+|S^{\prime}|-|S\cup S^{\prime}|\geq 1$, and we can choose $i=j\in S\cap S^{\prime}$. Furthermore, we have the inequalities
$\max\left\{{\left\lVert\sigma(S,k)\right\rVert}_{\infty},{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty}\right\}$ $\leq\sqrt{\Sigma^{\downarrow}_{1,1}}\,,$
and
$\max\left\{\sigma^{*}(S,k),\sigma^{*}(S^{\prime},k)\right\}$ $\leq\max_{i\in[d]}\sqrt{\Sigma^{\downarrow}_{i+k,i+k}\log(i+1)}\,,$
using which we deduce
$\Delta_{m}$ $\leq 2C\max_{i\in[d-k]}\sqrt{\Sigma^{\downarrow}_{i+k,i+k}\log(i+1)}+(2+2\sqrt{2})\beta_{r,m,k}\sqrt{\Sigma^{\downarrow}_{1,1}}\,.$
With $\Theta$ defined as in Proposition 3 we have
$\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}$ $=\left(\sum_{i=1}^{m}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}\kappa\,,$
and
$\sup_{\eta\in\Theta}\eta^{\downarrow}_{d-m+1}=\left(\sum_{i=1}^{d-m+1}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}\kappa\,.$
Therefore, Proposition 3 implies that any confidence interval for $\theta^{\downarrow}_{m}$ with coverage probability no less than $1-1/(2e)$, should have a width equal to $C_{\kappa}\left(\sum_{i=1}^{m}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}$ for some constant $C_{\kappa}\geq 0$ that may depend on $\kappa$. In particular, for any $\theta\in\Theta$ we have
$\operatorname{\mathbb{P}}\left(\Delta_{m}>C_{\kappa}\left(\sum_{i=1}^{m}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}\right)$ $\geq\frac{1}{2e}\,.$
Choosing $r=1+\log(8)\approx 3$, we have also shown that
$\operatorname{\mathbb{P}}\left(\Delta_{m}>2C\max_{i\in[d-k]}\sqrt{\Sigma^{\downarrow}_{i+k,i+k}\log(i+1)}+(2+2\sqrt{2})\beta_{r,m,k}\sqrt{\Sigma^{\downarrow}_{1,1}}\right)$ $\leq\frac{1}{2e}\,.$
Then, if we define
$p_{m}(\Sigma)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\sum_{i=1}^{m}\frac{\Sigma^{\downarrow}_{1,1}}{\Sigma^{\downarrow}_{i,i}}\,,$
and
$q_{m,k}(\Sigma)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\max_{i\in[d-k]}\frac{\Sigma^{\downarrow}_{i+k,i+k}}{\Sigma^{\downarrow}_{1,1}}\log(i+1)\,,$
then $\Delta_{m}$ is optimal up to a factor $\mathrm{polylog}(d)$, if $m$, $p_{m}(\Sigma)$, and $q_{m,k}(\Sigma)$ all bounded from above as $\mathrm{polylog}(d)$. Specifically, if $m$, $p_{m}(\Sigma)$, and $q_{m,k}(\Sigma)$ are all absolute constants, then $\Delta_{m}$ is optimal up to a constant factor.